OpenStudy (anonymous):

Solve the given equation: B - 4[B - 3(1 - B)] = 72

OpenStudy (midhun.madhu1987):

Open the innermost bracket by multiplying (1-B) with -3

OpenStudy (anonymous):

B - 4[B - 3 + B] = 72

OpenStudy (midhun.madhu1987):

Correct.. similarly second bracket also..

OpenStudy (anonymous):

B - 4B + 12 - 4B = 72

OpenStudy (midhun.madhu1987):

YEs... now combine the like terms and u get it...

OpenStudy (midhun.madhu1987):

wait....

OpenStudy (midhun.madhu1987):

B - 4[B - 3 + 3B] = 72 this is the first step ...u forgot to multiply -3 with B also.. Hope u understood the mistake

OpenStudy (anonymous):

will the answer be B = 64/-15 ?

OpenStudy (midhun.madhu1987):

nope.. can u check it again?

OpenStudy (anonymous):

-4?

OpenStudy (midhun.madhu1987):

Good... u did it.. :)

OpenStudy (anonymous):

thank you